OpenStudy (anonymous):

The product is 192 and the sum is minimum.

OpenStudy (anonymous):

at least what? The lowest value it can be?

OpenStudy (mary.rojas):

so yeah the lowest value..... well start off by listing all the possible pairs that have a product of 192

OpenStudy (mary.rojas):

sooo... like 12*16 and continue on

OpenStudy (anonymous):

I would break it down into prime numbers, and then use the entire list of prime numbers and multiply them together to make two numbers that make up 192. Then after a few you should have all of the numbers that could make up in a chart, and you can pick the two that add up to make the lowest number.

OpenStudy (anonymous):

x times y =192\[\rightarrow\] y=192/x x+y=s \[\rightarrow\] x+(192/x)=s ,, and then dervive both sides when you know s is a constant you get x=\[\sqrt{192}\] is a local and absolute minimum

OpenStudy (mary.rojas):

lol this question is from a week ago

OpenStudy (anonymous):

how embarrassing :S